# Spin Pumping into two-dimensional systems

### v1

@@ 4.1 Model
Here, $c_{{\bm{k}}\sigma}$ is an annihilation operator of conduction electrons with a wavenumber ${\bm{k}}=(k_{x},k_{y})$ and a spin $\sigma$ (=,absent=$\uparrow$,$\downarrow$= ↑ , ↓), $\xi_{\bm{k}}=\hbar^{2}{\bm{k}}^{2}/2m^{*}-\mu$ is the kinetic energy measured from the chemical potential $\mu$, $m^{*}$ is the effective mass, $\hat{I}$ is a $2\times 2$ identity matrix, and $\hat{\bm{\sigma}}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ denotes the Pauli matrices. The strengths of the Rashba and Dresselhaus spin-orbit interactions are denoted by $\alpha$ and $\beta$, respectively. The $2\times 2$ matrix $\hat{h}_{\bm{k}}$ can be rewritten as $\hat{h}_{\bm{k}}=\xi_{\bm{k}}\hat{I}-{\bm{h}}_{\rm eff}({\bm{k}})\cdot\hat{\bm{\sigma}}$, where ${\bm{h}}_{\rm eff}({\bm{k}})$ is an effective Zeeman field acting on the conduction electrons, which is defined as ${\bm{h}}_{\rm eff}(\varphi)\simeq k_{\rm F}(-\alpha\sin\varphi-\beta\cos\varphi,\alpha\cos\varphi+\beta\sin\varphi)$ using the polar representation $\hat{\bm{k}}=(\cos\varphi,\sin\varphi)$. The quantization axis of the electron spin depends on the electron wavenumber. As an illustrative example, we show the spin polarization near the Fermi surface for $\alpha/\beta=2$ in Fig. 6(b). The Hamiltonian for impurity scattering due to a short-range potential $v({\bm{r}})=v_{0}\delta({\bm{r}})$ is given as
@@
We introduce the proximity-induced exchange coupling between the FI and 2DEG explained in Sec. 2.3. We employ the model for a clean interface (see Sec. 2.3) [^5] It can be shown that the spin injection rate does not depend on the in-plain orientation of the magnetic field for an dirty interface.. Furthermore, we assume that the spin in FI points in the $xy$ plane and define a new coordinate $x^{\prime}y^{\prime}$ to make the $x^{\prime}$ direction align with the spin direction as shown in Fig. 6(c). Using the Holstein-Primakoff transformation explained in Sec. 2.3, the Hamiltonian for the proximity-induced exchange coupling is given as
@@ 4.2 Enhancement of Gilbert damping
4.3 Inverse Rashba-Edelstein effect
@@
where $E^{\gamma}_{\bm{k}}$ is an energy dispersion of electrons in 2DEG, $f_{0}(\epsilon)=(\exp[(\epsilon-\mu)/k_{\rm B}T]+1)^{-1}$ is the Fermi distribution function, and $\delta\mu(\varphi,\gamma)$ is a chemical potential shift in the direction of $\varphi$. Combining these expressions of the collision terms with the Boltzmann equation, we can calculate the nonequilibrium distribution function $f({\bm{k}},\gamma)$ (for a detailed calculation, see Ref. 116). Note that our formulation does not need the concept of spin current.
@@
The two contour plots on the left of Fig. 8 show the current density as functions of the FMR frequency and the azimuth angle of the spin in FI for $\alpha/\beta=3$. Due to the distribution of $h_{\rm eff}(\varphi)$, the current densities exhibit large values across a wide range of $4k_{\rm F}\beta\lesssim\hbar\omega_{\bm{0}}\lesssim 8k_{\rm F}\beta$. Near the lower boundary at $\hbar\omega_{\bm{0}}/k_{\rm F}\beta=4$, the current amplitude vanishes at $\theta=3\pi/4$ (point A in the contour plot) and takes a maximum at $\theta=\pi/4$ (point B), where the current flows along the $(1,-1)$ direction. On the other hand, near the higher boundary at $\hbar\omega_{\bm{0}}/k_{\rm F}\beta=8$, the current amplitude takes a maximum at $\theta=3\pi/4$ (point C), where the current flows along the $(1,1)$ direction, and vanishes at $\theta=\pi/4$ (point D). The Fermi surface modulations at the corresponding points are schematically shown in the right four pictures of Fig. 8. It is remarkable that the full solution of the Boltzmann equation is needed for an accurate description of the IREE. If we employ the relaxation-time approximation [145, 146], the results are changed significantly as shown in the middle contour plots of Fig. 8.
@@ 5 Spin pumping into TMDC
Figure 9: Distinctive properties of monolayer TMDCs. (a) When viewed from above, the lattice structure forms a honeycomb lattice. The unit cell lacks a inversion center, breaking spatial inversion symmetry. (b) Schematic picture of the band structure of monolayer TMDCs. A Zeeman-type spin splitting occurs due to the strong SOC of the transition metals. Because of the mirror symmetry of the crystal structure, the $z$-component of the spin becomes a conserved quantity up to first order of wave number in the model Hamiltonian. (c) Schematic picture of valley-dependent Berry curvature. The Berry curvature exhibits peaks in both $K_{+}$ and $K_{-}$ valleys, with their signs being opposite.
@@
In this section, we theoretically investigate valley-dependent spin transport phenomena induced by spin excitation in a monolayer TMDC. Figure 10 shows a schematic of the system, where an FI is attached to a monolayer TMDC. Microwave irradiation is then applied to the system, triggering FMR. Through spin-transfer processes mediated by proximity exchange coupling at the interface, the FMR excites electron spins in the monolayer TMDC. Our findings reveal that the interplay between SVC and proximity exchange coupling leads to valley-selective spin excitation. Owing to the valley-dependent Berry curvature, this excitation gives rise to transverse spin accumulation—a phenomenon we term the spin-current Hall effect.
Figure 10: A schematic illustration of the system where the FI is attached to the TMDC. Under microwave irradiation, spin pumping occurs at the interface between the FI and the TMDC, leading to the generation of a diffusive spin current in the TMDC.
@@ 5.2 Interface spin current
where we introduce the valley-resolved spin current:
@@
[…]
@@ 5.3 Numerical results
A fundamental result derived from the above expressions is the feasibility of generating a valley-polarized spin current. This is because the valley degeneracy is lifted by the proximity exchange coupling, allowing the spin current in each valley to differ. To characterize the valley polarization of the spin current, we define the valley-polarized spin current (VPSC) as $I_{\mathrm{s}}^{K_{+}}-I_{\mathrm{s}}^{K_{-}}$. We show that, with appropriate carrier doping, spins can be valley-selectively excited, resulting in a completely valley-polarized spin current. Panels (a) and (b) of Fig. 11 show the valence bands in the $K_{+}$ and $K_{-}$ valleys, respectively. We set the parameters as $\lambda/\Delta=0.10$ and $JS_{0}/\Delta=0.05$, consistent with values from first-principles calculations [176, 177, 178, 179]. Panels (c), (d), and (e) of Fig. 11 show the spin current in the $K_{+}$ valley, the spin current in the $K_{-}$ valley, and the VPSC, respectively. In energy region (i), the spin current at zero temperature is finite only in the $K_{+}$ valley, indicating valley-selective spin excitation and the generation of a fully valley-polarized spin current. This condition remains valid even at finite temperatures provided that the spin splitting due to proximity exchange coupling is much greater than thermal broadening, i.e., $JS_{0}\gg k_{\mathrm{B}}T$. In energy region (ii), however, the spin current is finite in both valleys, leading to an almost zero VPSC and suppressed valley selectivity. We also note that a small spin splitting exists in the conduction band [169, 170, 171, 172], which is omitted in our model Hamiltonian. Thus, valley-selective spin excitation is also feasible in the conduction band.
The spin current at the interface induces spin accumulation in the TMDC, which in turn drives a diffusive spin current within the TMDC. This diffusive spin current is schematically shown in Fig. 11(f), and consists of the flow of up-spin electrons and the counterflow of down-spin electrons. In energy region (i), the diffusive spin current originates from electrons in the $K_{+}$ valley. The Berry curvature has the same sign for both up-spin and down-spin electrons within the same valley, giving rise to an intrinsic Hall conductivity [153, 180, 23]. Thus, the diffusive spin current results in a transverse spin accumulation—a phenomenon we refer to as the ”spin-current Hall effect”. In contrast, in energy region (ii), the diffusive spin current arises from electrons in both the $K_{+}$ and $K_{-}$ valleys. Since these valleys are related by time-reversal symmetry, their Hall conductivities have opposite signs. As a result, the transverse spin accumulations from the $K_{+}$ and $K_{-}$ valleys cancel each other.
@@
Figure 11: The panels (a) and (b) show the valence bands in the $K_{+}$ and $K_{-}$ valleys, respectively. The panels (c) and (d) show the spin current at several temperatures in the $K_{+}$ and $K_{-}$ valleys, respectively. The panel (e) shows the valley-polarized spin current at several temperatures. The units of the spin current are given by $I_{0}=\frac{|J|^{2}\mathcal{A}(S_{0}\gamma_{\mathrm{g}}h_{\mathrm{ac}})^{2}\hbar\omega}{(\omega-\omega_{\bm{0}})^{2}+\alpha_{\mathrm{G}}^{2}\omega_{\bm{0}}^{2}}\frac{\Delta^{2}}{(\hbar v)^{4}}$. (f) (i) The up-spin (blue arrows) and down-spin (red arrows) electrons flowing in opposite directions lead to the transverse spin accumulation. (ii) When the spin current is finite in the $K_{+}$ and $K_{-}$ valleys, the transverse spin accumulation almost cancels.
Finally, we turn our attention to a potential experimental technique for detecting transverse spin accumulation. The magneto-optical Kerr effect offers a promising approach for identifying such accumulation [181, 182, 164]. High-resolution imaging of the Kerr angle can reveal the spatial distribution of the spin accumulation. Despite strong spin–orbit coupling in monolayer TMDCs, the out-of-plane spin diffusion length is expected to be long enough to be resolved by Kerr microscopy, whose spatial resolution is about one micron.
@@ 6 Other Applications
Orbital current, i.e., the flow of orbital angular momentum, has attracted considerable attention in the field of spintronics and has been extensively studied [191, 192, 193]. Among the various phenomena associated with orbital current, orbital pumping, which is an extension of spin pumping to orbital angular momentum, has been investigated both experimentally and theoretically [194, 195]. A detailed investigation of orbital pumping based on microscopic models remains an important direction for future research.
@@ 7 Summary
Finally, we examined spin transport phenomena in monolayer transition metal dichalcogenides. These materials, characterized by strong spin–orbit coupling and broken inversion symmetry, exhibit spin–valley coupling and finite Berry curvature effects. We have shown that spin pumping into monolayer transition metal dichalcogenides can induce valley-selective spin excitations, resulting in transverse spin accumulation under the injection of a longitudinal spin current. This mechanism gives rise to what we term the spin-current Hall effect, offering new routes for controlling valley degrees of freedom via spin excitations.
@@ Acknowledgments
The authors are grateful to J. Fujimoto, T. Funato, H. Funaki, K. Hosokawa, M. Kohda, Y. Suzuki, and Y. Kato for valuable discussions. This work was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12374126, by the Priority Program of Chinese Academy of Sciences under Grant No. XDB28000000, and by JSPS KAKENHI for Grants (Nos. JP21H01800, JP21H04565, JP23H01839, JP24K06951, JP24H00322, JP24H00853, and JP25K07224) from MEXT, Japan.
Appendix A Holstein Primakoff transformation
@@
where we require $[b_{n},b_{m}^{\dagger}]=\delta_{n,m}$ to ensure that $S_{n}^{+}$, $S_{n}^{-}$, and $S_{n}^{z}$ satisfy the commutation relation of angular momentum. The deviation of $\langle S_{n}^{z}\rangle$ from its maximum value $S_{0}$ is quantified using the boson particle number $\langle b_{n}^{\dagger}b_{n}\rangle$. Assuming a small deviation such that $\langle b_{n}^{\dagger}b_{n}\rangle/S_{0}\ll 1$, the ladder operators $S_{n}^{\pm}$ can be approximated by $S_{n}^{+}\simeq(2S_{0})^{1/2}b_{n}$ and $S_{n}^{-}\simeq(2S_{0})^{1/2}b_{n}^{\dagger}$. This is known as the spin-wave approximation. Performing Fourier expansion, the boson operators are given by
@@
where $N_{\rm FI}$ is the number of sites. The boson operators $b_{{\bf\it k}}^{\dagger}$ and $b_{{\bf\it k}}$ with wave vector ${\bf\it k}$ satisfy $[b_{{\bf\it k}},b_{{\bf\it k}^{\prime}}^{\dagger}]=\delta_{{\bf\it k},{\bf\it k}^{\prime}}$ and describe the collective excitation called magnon. Substituting these approximate relations, the Hamiltonian of the FI, Eq. (7), is modified as Eq. (9) in the leading order of $1/S_{0}$.

### v2

@@ 4.1 Model
Here, $c_{{\bm{k}}\sigma}$ is an annihilation operator of conduction electrons with a wavenumber ${\bm{k}}=(k_{x},k_{y})$ and a spin $\sigma$ ($\uparrow$$\downarrow$), $\xi_{\bm{k}}=\hbar^{2}{\bm{k}}^{2}/2m^{*}-\mu$ is the kinetic energy measured from the chemical potential $\mu$, $m^{*}$ is the effective mass, $\hat{I}$ is a $2\times 2$ identity matrix, and $\hat{\bm{\sigma}}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ denotes the Pauli matrices. The strengths of the Rashba and Dresselhaus spin-orbit interactions are denoted by $\alpha$ and $\beta$, respectively. The $2\times 2$ matrix $\hat{h}_{\bm{k}}$ can be rewritten as $\hat{h}_{\bm{k}}=\xi_{\bm{k}}\hat{I}-{\bm{h}}_{\rm eff}({\bm{k}})\cdot\hat{\bm{\sigma}}$, where ${\bm{h}}_{\rm eff}({\bm{k}})$ is an effective Zeeman field acting on the conduction electrons, which is defined as ${\bm{h}}_{\rm eff}(\varphi)\simeq k_{\rm F}(-\alpha\sin\varphi-\beta\cos\varphi,\alpha\cos\varphi+\beta\sin\varphi)$ using the polar representation $\hat{\bm{k}}=(\cos\varphi,\sin\varphi)$. The quantization axis of the electron spin depends on the electron wavenumber. As an illustrative example, we show the spin polarization near the Fermi surface for $\alpha/\beta=2$ in Fig. 6(b). The Hamiltonian for impurity scattering due to a short-range potential $v({\bm{r}})=v_{0}\delta({\bm{r}})$ is given as
@@
We introduce the proximity-induced exchange coupling between the FI and 2DEG explained in Sec. 2.3. We employ the model for a clean interface (see Sec. 2.3) [^5] It can be shown that the spin injection rate does not depend on the in-plane orientation of the magnetic field for a dirty interface.. Furthermore, we assume that the spin in FI points in the $xy$ plane and define a new coordinate $x^{\prime}y^{\prime}$ to make the $x^{\prime}$ direction align with the spin direction as shown in Fig. 6(c). Using the Holstein-Primakoff transformation explained in Sec. 2.3, the Hamiltonian for the proximity-induced exchange coupling is given as
@@ 4.2 Enhancement of Gilbert damping
4.3 Inverse Rashba-Edelstein effect (IREE)
@@
where $E^{\gamma}_{\bm{k}}$ is an energy dispersion of electrons in 2DEG, $f_{0}(\epsilon)=(\exp[(\epsilon-\mu)/k_{\rm B}T]+1)^{-1}$ is the Fermi distribution function, and $\delta\mu(\varphi,\gamma)$ is a chemical potential shift in the direction of $\varphi$. Combining these expressions of the collision terms with the Boltzmann equation, we can calculate the nonequilibrium distribution function $f({\bm{k}},\gamma)$ (for a detailed calculation, see Ref. 126). Note that our formulation does not need the concept of spin current.
@@
The two contour plots on the left of Fig. 8 show the current density as functions of the FMR frequency and the azimuth angle of the spin in FI for $\alpha/\beta=3$. Due to the distribution of $h_{\rm eff}(\varphi)$, the current densities exhibit large values across a wide range of $4k_{\rm F}\beta\lesssim\hbar\omega_{\bm{0}}\lesssim 8k_{\rm F}\beta$. Near the lower boundary at $\hbar\omega_{\bm{0}}/k_{\rm F}\beta=4$, the current amplitude vanishes at $\theta=3\pi/4$ (point A in the contour plot) and takes a maximum at $\theta=\pi/4$ (point B), where the current flows along the (1,1)(1,-1) direction. On the other hand, near the higher boundary at $\hbar\omega_{\bm{0}}/k_{\rm F}\beta=8$, the current amplitude takes a maximum at $\theta=3\pi/4$ (point C), where the current flows along the $(1,1)$ direction, and vanishes at $\theta=\pi/4$ (point D). The Fermi surface modulations at the corresponding points are schematically shown in the right four pictures of Fig. 8. It is remarkable that the full solution of the Boltzmann equation is needed for an accurate description of the IREE. If we employ the relaxation-time approximation [153, 154], the results are changed significantly as shown in the middle contour plots of Fig. 8.
@@ 5 Spin pumping into TMDC
Figure 9: Distinctive properties of monolayer TMDCs. (a) When viewed from above, the lattice structure forms a honeycomb lattice. The unit cell lacks a inversion center, breaking spatial inversion symmetry. (b) Schematic picture of the band structure of monolayer TMDCs. A Zeeman-type spin splitting occurs due to the strong SOC of the transition metals. Because of the mirror symmetry of the crystal structure, the $z$-component of the spin becomes a conserved quantity up to first order of wave number in the model Hamiltonian. (c) Schematic picture of valley-dependent Berry curvature. The Berry curvature exhibits peaks in both $K_{+}$ and $K_{-}$ valleys, with their signs being opposite. (d) A schematic illustration of the system where the FI is attached to the TMDC. Under microwave irradiation, spin pumping occurs at the interface between the FI and the TMDC.
@@
In this section, we theoretically investigate the spin pumping into a monolayer TMDC [177]. Figure 9 (d) shows a schematic of the system, where an FI is attached to a monolayer TMDC. Microwave irradiation is then applied to the system, triggering FMR. Through spin-transfer processes mediated by proximity exchange coupling at the interface, the FMR excites electron spins in the monolayer TMDC. Our findings reveal that the interplay between SVC and proximity exchange coupling leads to valley-selective spin excitation.
@@ 5.2 Interface spin current
where we introduce the valley-resolved spin current
@@
$\sum_{{\bf\it q}}\mathrm{Im}\,\chi^{R}_{\tau}({\bf\it q},\omega)=-2\pi\hbar\omega\int d\varepsilon\left(-\frac{\partial f(\varepsilon)}{\partial\varepsilon}\right)D_{\tau,+}(\varepsilon)D_{\tau,-}(\varepsilon)\left(1+\frac{Z_{\tau,+}Z_{\tau,-}}{|\varepsilon-E_{\tau,+}|\,|\varepsilon-E_{\tau,-}|}\right),$ (71)
@@ 5.3 Numerical results
A fundamental result derived from the above expressions is the feasibility of generating a valley-polarized spin current (VPSC). This is because the valley degeneracy is lifted by the proximity exchange coupling, allowing the spin current in each valley to differ. To characterize the valley polarization of the spin current, we define the VPSC as $I_{\mathrm{s}}^{K_{+}}-I_{\mathrm{s}}^{K_{-}}$. We show that, with appropriate carrier doping, spins can be valley-selectively excited, resulting in a completely VPSC. Panels (a) and (b) of Fig. 10 show the valence bands in the $K_{+}$ and $K_{-}$ valleys, respectively. We set the parameters as $\lambda/\Delta=0.10$ and $JS_{0}/\Delta=0.05$, consistent with values from first-principles calculations [185, 186, 187, 188]. Panels (c), (d), and (e) of Fig. 10 show the spin current in the $K_{+}$ valley, the spin current in the $K_{-}$ valley, and the VPSC, respectively. In energy region (i), the spin current at zero temperature is finite only in the $K_{+}$ valley, indicating valley-selective spin excitation and the generation of a completely VPSC. This condition remains valid even at finite temperatures provided that the spin splitting due to proximity exchange coupling is much greater than thermal broadening, i.e., $JS_{0}\gg k_{\mathrm{B}}T$. In energy region (ii), however, the spin current is finite in both valleys, leading to an almost zero VPSC and suppressed valley selectivity. We also note that a small spin splitting exists in the conduction band [178, 179, 180, 181], which is omitted in our model Hamiltonian. Thus, valley-selective spin excitation is also feasible in the conduction band.
The present review concentrates on monolayer TMDCs, while extending the theoretical framework to multilayers remains an open and important problem. Spin pumping into multilayer TMDCs has already been demonstrated experimentally [71, 75], highlighting the significance of this research avenue. As the layer number increases, the band gap evolves from direct to indirect [20, 160]. Symmetry analysis indicates that inversion symmetry is absent in odd layers, which lifts spin degeneracy and enables spin–valley coupling, whereas even layers preserve inversion symmetry and remain spin-degenerate [21]. Odd-layer systems are therefore expected to exhibit physics analogous to that discussed in this review, and a detailed theoretical investigation will be pursued in future work.
@@
Figure 10: The panels (a) and (b) show the valence bands in the $K_{+}$ and $K_{-}$ valleys, respectively. The panels (c) and (d) show the spin current at several temperatures in the $K_{+}$ and $K_{-}$ valleys, respectively. The panel (e) shows the valley-polarized spin current at several temperatures. The units of the spin current are given by $I_{0}=\frac{|J|^{2}\mathcal{A}(S_{0}\gamma_{\mathrm{g}}h_{\mathrm{ac}})^{2}\hbar\omega}{(\omega-\omega_{\bm{0}})^{2}+\alpha_{\mathrm{G}}^{2}\omega_{\bm{0}}^{2}}\frac{\Delta^{2}}{(\hbar v)^{4}}$.
@@ 6 Other Applications
Spin pumping has also been extended to organic semiconductors, attractive materials due to their potential for long spin lifetimes and chemical tunability [197]. A pioneering demonstration of spin-charge conversion via the ISHE was achieved in a solution-processed conducting polymer [198]. A key advantage of these molecular systems is the ability to engineer their properties. Indeed, it was subsequently shown that the ISHE can be controlled by chemically tuning the spin-orbit coupling of the molecules [199]. The potential of this approach has been further underscored by reports of giant spin pumping at an optimized ferromagnet-organic interface, with efficiencies comparable to those of heavy metals [200]. However, reliably quantifying the ISHE remains a central challenge, as the genuine signal is often obscured by parasitic thermoelectric and spin rectification effects, necessitating careful experimental methodologies to ensure accurate interpretation [201, 202].
Recently, orbital current, i.e., the flow of orbital angular momentum, has attracted considerable attention in the field of spintronics and has been extensively studied [203, 204, 205]. Among the various phenomena associated with orbital current, orbital pumping, which is an extension of spin pumping to orbital angular momentum, has been investigated both experimentally and theoretically [206, 207]. A detailed investigation of orbital pumping based on microscopic models remains an important direction for future research.
@@ 7 Summary
Finally, we examined spin transport phenomena in monolayer transition metal dichalcogenides (TMDCs). These materials, characterized by strong spin–orbit coupling and broken inversion symmetry, exhibit spin–valley coupling and finite Berry curvature effects. We have shown that spin pumping into these TMDCs can induce valley-selective spin excitations, offering new routes for controlling valley degrees of freedom via spin excitations.
@@ Acknowledgments
The authors are grateful to J. Fujimoto, T. Funato, H. Funaki, K. Hosokawa, M. Kohda, Y. Suzuki, and Y. Kato for valuable discussions. This work was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12374126, by the Priority Program of Chinese Academy of Sciences under Grant No. XDB28000000, Waseda University Grant for Special Research Projects (Grants No. 2025C-651 and No. 2025R-061), and by JSPS KAKENHI for Grants (Nos. JP21H01800, JP21H04565, JP23H01839, JP24K06951, JP24H00322, JP24H00853, JP24KJ0624, and JP25K07224) from MEXT, Japan.
Appendix A Holstein-Primakoff transformation
@@
where we require $b_{n},b_{m}^{\dagger}]=\delta_{n,m}$ to ensure that $S_{n}^{+}$, $S_{n}^{-}$, and $S_{n}^{z}$ satisfy the commutation relation of angular momentum. The deviation of $\langle S_{n}^{z}\rangle$ from its maximum value $S_{0}$ is quantified using the boson particle number $\langle b_{n}^{\dagger}b_{n}\rangle$. Assuming a small deviation such that $\langle b_{n}^{\dagger}b_{n}\rangle/S_{0}\ll 1$, the ladder operators $S_{n}^{\pm}$ can be approximated by $S_{n}^{+}\simeq(2S_{0})^{1/2}b_{n}$ and $S_{n}^{-}\simeq(2S_{0})^{1/2}b_{n}^{\dagger}$. This is known as the spin-wave approximation. Performing Fourier expansion, the boson operators are given by
@@
where $N_{\rm FI}$ is the number of sites. The boson operators $b_{{\bf\it k}}^{\dagger}$ and $b_{{\bf\it k}}$ with wave vector $\bf\it k$} satisfy $b_{{\bf\it k}},b_{{\bf\it k}^{\prime}}^{\dagger}]=\delta_{{\bf\it k},{\bf\it k}^{\prime}}$ and describe the collective excitation called magnon. Substituting these approximate relations, the Hamiltonian of the FI, Eq. (7), is modified as Eq. (9) in the leading order of $1/S_{0}$.